30th July (Thursday)

  • Data Processing: data "stripping" productions ongoing. User and MC jobs

  • T0
    • Discussion with LSF team about wall and cpu time queue lenghts, ongoing (GGUS:115027)

  • T1
    • RAL: LAN network instabilities, disk storages were automatically banned on Wednesday
    • SARA: unscheduled outage of CEs

27th July (Monday)

  • Data Processing: data "stripping" productions ongoing. User and MC jobs

  • T0
    • Discussion with LSF team about wall and cpu time queue lenghts, ongoing (GGUS:115027)

  • T1
    • IN2P3: getting low on user disk space, some more disk to be installed

23th July (Thursday)

  • Data Processing: "new data" stripping. User and MC jobs

  • T0
    • NTR

  • T2 * NTR

20th July (Monday)

  • Data Processing: "new data" automatic production in place. User and MC jobs

  • T0
    • NTR

  • T1
    • NTR

  • T2 * NTR

16th July (Thursday)

Half of the jobs having problem to contact a vobox at CERN. Investigating.

9th July (Thursday)

  • Data Processing, User and MC jobs

  • T0
    • NTR

6th July (Monday)

  • Data Processing, User and MC jobs on the grid.
  • T0
    • Data transfer problem from Pit to CASTOR. Fixed, misconfiguration of some new SRM nodes.

2nd July (Thursday)

  • Validating Data Processing workflows. User and MC jobs on the grid.
  • T0
  • T1

-- JoelClosier - 2016-01-07

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2016-01-07 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback